1.We purchased a building with a market value of $150,000 for $125,000 cash. The seller paid $50,000 when they originally purchased the building. What amount is used when we record the purchase of the building?

$25,000

$50,000

$125,000

$150,000

2.On July 1, 2017, we purchased a truck for $60,000. The truck has a useful life of 5 years and a residual value of $5,000. The truck is depreciated using the straight-line method. What is the depreciation expense for 2017?

$12,000

$11,000

$5,500

$1,000

3.On July 1, 2017, we purchased a truck for $60,000. The truck has a useful life of 5 years and a residual value of $5,000. The truck is depreciated using the double-declining balance method. What is the depreciation expense for 2018 (Year 2)?

$12,000

$19,200

$13,200

$11,000

4.On January 1, 2017, we purchased a truck for $55,000. The truck has a useful life of 5 years or 100,000 miles, and a residual value of $5,000. The actual mileage for the truck was 25,000 miles in 2017 and 30,000 miles in 2018. The truck is depreciated using the units of production method. What is the balance in accumulated depreciation expense after the adjusting entry for 2017?

$12,500

$15,000

$22,000

$20,000

5.A truck cost $60,000 when purchased and has a $55,000 balance in the accumulated depreciation account. If the truck is discarded we will record:

loss on disposal, $5,000

gain on disposal, $5,000

no gain or loss because it was discarded

loss of $60,000
Answer #1
1
Amount recorded as purchase of the building 125000
Option 3 $125,000 is correct
2
Cost of truck 60000
Less:Residual value 5000
55000
Divide by useful life 5
Annual depreciation 11000
Fraction of year in 2017 6/12
Depreciation expense for 2017 5500
Option 3 $5,000 is correct
3
Double-declining balance rate 40% =1/5*2
Depreciation expense for 2017 12000 =60000*40%*6/12
Depreciation expense for 2018 19200 =(60000-12000)*40%
Option 2 $19,200 is correct
4
Cost of truck 55000
Less:Residual value 5000
50000
Divide by Total miles 100000
Depreciation per mile 0.5
X Actual mileage for 2017 25000
Depreciation expense for 2017 12500
Balance in accumulated depreciation 12500
Option 1 $12,500 is correct
5
Loss on disposal, $5,000 (60000-55000)
Option 1 is correct